# Bandits with Abstention under Expert Advice

### v1

@@ Abstract
We study the classic problem of prediction with expert advice under bandit feedback. Our model assumes that one action, corresponding to the learner’s abstention from play, has no reward or loss on every trial. We propose the CBA algorithm, which exploits this assumption to obtain reward bounds that can significantly improve those of the classical Exp4 algorithm. We can view our problem as the aggregation of confidence-rated predictors when the learner has the option of abstention from play. Importantly, we are the first to achieve bounds on the expected cumulative reward for general confidence-rated predictors. In the special case of specialists we achieve a novel reward bound, significantly improving previous bounds of SpecialistExp (treating abstention as another action). As an example application, we discuss learning unions of balls in a finite metric space. In this contextual setting, we devise an efficient implementation of CBA, reducing the runtime from quadratic to almost linear in the number of contexts. Preliminary experiments show that CBA improves over existing bandit algorithms.
Multi Armed Bandits, Online Learning
@@ 1 Introduction
We study the classic problem of prediction with expert advice under bandit feedback. The problem is structured as a sequence of trials. During each trial, each expert recommends a probability distribution over the set of possible actions. The learner then selects an action, observes, and incurs the (potentially negative) reward associated with that action on that particular trial. In practical applications, errors often lead to severe consequences, and consistently making predictions is neither safe nor economically practical. For this reason, the abstention option has gained a lot of interest in the literature, both in the batch and online setting (Chow, 1957, 1970; Hendrickx et al., 2021; Cortes et al., 2018). Similarly to previous works, this paper is based on the assumption that one of the actions always has zero reward: such an action is equivalent to an abstention of the learner from play. Apart from the fact that the rewards are bounded, we make no further assumptions on how the rewards or expert predictions are generated. In this paper, we present an efficient algorithm CBA (Confidence-rated Bandits with Abstentions) which exploits the abstention action to get reward bounds that can be dramatically higher than those of Exp4 (Auer et al., 2002). In the worst case, our reward bound essentially matches that of Exp4 so that CBA can be seen as a strict improvement, since the time-complexities of the two algorithms are identical in the general case.
@@
Our problem can also be seen as that of aggregating confidence-rated predictors (Blum & Mansour, 2007; Gaillard et al., 2014; Luo & Schapire, 2015) when the learner has the option of abstaining from taking actions. When the problem is phrased in this way, at the start of each trial, each predictor recommends a probability distribution over the actions (which now may not include an action with zero reward) but with a confidence rating. A low confidence rating can mean that either the predictor thinks that all actions are bad (so that the learner should abstain) or simply does not know which action is the best. Previous works measure the performance of their algorithm in terms of a sum of scaled per-trial rewards.
In contrast to previous algorithms for confidence-rated predictors, our approach allows for the derivation of bounds on the expected cumulative reward of CBA. One specific scenario where prior algorithms can establish cumulative reward bounds is when the predictors are specialists (Freund et al., 1997), having either full confidence (a.k.a. awake) or no confidence (a.k.a. asleep) on any given trial. The SpecialistExp algorithm, a bandit version of the standard specialist algorithm (Herbster et al., 2021), achieves regret bounds with respect to any subset of specialists in which exactly one specialist is awake on each trial. In Section 6, we present an illustrative problem involving learning balls in a space equipped with a distance function. This example demonstrates our capability to significantly outperform SpecialistExp, which treats abstention as just another action, when our confidence-rated predictors are indeed specialists. For this problem, we will give subroutines that significantly speed up CBA. We note that, for any specialist problem, in the worst case the cumulative reward of CBA is essentially equal to that of SpecialistExp so that it can be seen as a strict improvement.
@@ 1.1 Additional Related Work
The non-stochastic multi-armed bandit problem, initially introduced by Auer et al. (2002), has been a subject of significant research and interest. Auer et al. (2002) also introduced the multi-armed bandit problem with expert advice, introducing the Exp4 algorithm. The Exp4 algorithm, evolved the field of traditional multi-armed bandits to encompass more complex scenarios, particularly the contextual bandit setting (Lattimore & Szepesvári, 2020). Contextual bandits represent an extension of the classical multi-armed bandit framework, where an agent makes a sequence of decisions while taking into account contextual information.
As discussed in the introduction, a key aspect of this work is the option to abstain from making any decision. In the batch setting (Chow, 1957, 1970), this is usually referred to as machine learning with “rejection”, as these works study whether to utilize or reject a specific model prediction based on the specific request (see Hendrickx et al. (2021) for a survey). This line of research translated into the online learning world with the possibility of abstention by the learner. In these works, the authors usually rely on a cost associated with the abstention action. Neu & Zhivotovskiy (2020) studied the magnitude of the cost associated with abstention in an expert setting with bounded losses. They state that if the cost is lower than half of the amplitude of the interval of the loss, it is possible to derive bounds that are independent of the time horizon. In Cortes et al. (2018) a non-contextual and partial information setting with the option of abstention is studied. The sequel model (Cortes et al., 2020) regards this model as a special case of their stochastic feedback graph model. Schreuder & Chzhen (2021) also studied the possible implication in the fairness setting when utilizing the option of abstaining as it could lead to discriminatory predictions.
Our work is also related to the multi-class classification with bandit feedback introduced as weak reinforcement (Auer & Long, 1999). An action in our bandit setting corresponds to a class in the multi-class classification framework.
The study most akin to ours is that of Herbster et al. (2021), which introduced the SpecialistExp algorithm, a bandit variant of the standard Specialists algorithm from Freund et al. (1997). We differ from these works as the abstention is treated as an algorithmic choice, introducing a variant of the specialist framework. Instead of sleeping in the rounds where the specialist is not active, the specialist will vote for abstention, which is a proper action of our algorithm.
@@ 2 Problem Formulation and Notation
    The reward incurred by action $a_{t}$ on trial $t$ (which is in $[-1,1]$ and is selected by Nature before the trial) is revealed to the learner.
The aim of the learner is to maximize the cumulative reward obtained by its selected actions. As discussed in Section 1, we consider the case where there is an action that incurs zero reward on every trial, for the abstention action.
@@
Any vector $\boldsymbol{s}\in\mathcal{A}$ represents the probability distribution over actions which assigns, for all $a\in[K]$, a probability of $s_{a}$ to action $a$, and assigns a probability of $1-\|\boldsymbol{s}\|_{1}$ to the abstention action $\square$, where $\|\boldsymbol{s}\|_{1}$ denotes 1-norm of $\boldsymbol{s}$. We write $a\sim\boldsymbol{s}$ to represent that an action $a$ is drawn from this probability distribution. We will refer to the elements of the set $\mathcal{A}$ as stochastic actions.
A policy is any element of $\mathcal{A}^{T}$ (noting that any such policy is a matrix in $[0,1]^{T\times K}$). Any policy $\boldsymbol{e}\in\mathcal{A}^{T}$ defines a stochastic sequence of actions: on every trial $t\in[T]$ an action $a\in[K]\cup\{\square\}$ being drawn as $a\sim\boldsymbol{e}_{t}$. Note that if the learner plays according to a policy $\boldsymbol{e}\in\mathcal{A}^{T}$ then on each trial $t$ it obtains an expected reward of $\boldsymbol{r}_{t}\cdot\boldsymbol{e}_{t}$, where the operator $\cdot$ denotes the dot product. Note that each expert is equivalent to a policy. Thus, for all $i\in[E]$ we denote the $i$-th expert by $\boldsymbol{e}^{i}\in\mathcal{A}^{T}$. At the start of each trial $t\in[T]$, the learner views the sequence $\langle\boldsymbol{e}^{i}_{t}\,|\,i\in[E]\rangle$.
We can also view the experts alternatively as confidence-rated predictors over the set $[K]$: For each $i\in[E]$ and $t\in[T]$, the vector $\boldsymbol{e}^{i}_{t}$ can be viewed as suggesting the probability distribution $\boldsymbol{e}^{i}_{t}/\|\boldsymbol{e}^{i}_{t}\|_{1}$ over $[K]$, but with confidence $\|\boldsymbol{e}^{i}_{t}\|_{1}$. We denote this confidence by $c_{t,i}:=\|\boldsymbol{e}^{i}_{t}\|_{1}$ and write $\boldsymbol{c}_{t}:=(c_{t,1},\dots,c_{t,E})$.
In this work we will refer to the unnormalized relative entropy defined by:
$\Delta(\boldsymbol{u},\boldsymbol{v}):=\sum_{i\in[E]}u_{i}\ln\left(\frac{u_{i}}{v_{i}}\right)-\|\boldsymbol{u}\|_{1}+\|\boldsymbol{v}\|_{1},$
@@ 3 Main Result
Our main result is represented by the bound on the cumulative reward of our algorithm CBA. We note that any weight vector $\boldsymbol{u}\in\mathbb{R}_{+}^{E}$ induces a matrix $\boldsymbol{\pi}(\boldsymbol{u})\in\mathbb{R}_{+}^{T\times K}$ defined by
@@
which is the linear combination of the experts with coefficients given by $\boldsymbol{u}$. However, only some such linear combinations generate valid policies. Thus, we define
$\mathcal{V}:=\{\boldsymbol{u}\in\mathbb{R}_{+}^{E},|,\boldsymbol{\pi}(\boldsymbol{u})\in\mathcal{A}^{T}\}$
as the set of all weight vectors that generate valid policies. Particularly, note that $\boldsymbol{u}\in\mathcal{V}$ if and only if, on every trial $t$, the weighted sum of the confidences $\boldsymbol{u}\cdot\boldsymbol{c}_{t}$ is no greater than one. Given some $\boldsymbol{u}\in\mathcal{V}$ we define
@@
The following theorem (proved in Section 5) allows us to bound the regret of CBA with respect to any valid linear combination $\boldsymbol{u}$ of experts.
@@
$\sum_{t\in[T]}\mathbb{E}[r_{t,a_{t}}]\geq\mathbb{E}[\rho(\boldsymbol{u})]-\frac{\Delta(\boldsymbol{u},\boldsymbol{w}_{1})}{\eta}-\eta(12K+2)T$
where the expectations are with respect to the randomization of CBA’s strategy. The per-trial time runtime complexity of CBA is in $\mathcal{O}(KE)$.
We now compare our bound to that of previous algorithms. Firstly, Exp4 can only achieve bounds relative to a $\boldsymbol{u}\in\mathcal{V}$ with $\|\boldsymbol{u}\|_{1}=1$ , in which case it essentially matches our bound but with $212K+$ 2 replaced by $88K+$ 8. Hence, for any $\boldsymbol{u}\in\mathcal{V}$ the Exp4 bound essentially replaces the term $\rho(\boldsymbol{u})$ in our bound by $\rho(\boldsymbol{u})/\|\boldsymbol{u}\|_{1}$. Note that $\|\boldsymbol{u}\|_{1}$ could be as high as the number of experts which implies we can dramatically outperform Exp4. When viewing our experts as confidence-rated predictors, we note that previous algorithms for this setting only give bounds on a weighted sum of the per-trial rewards where the weight on each trial is $\boldsymbol{u}\cdot\boldsymbol{c}_{t}$ for some $\boldsymbol{u}\in\mathcal{V}$. This is only a cumulative reward bound when $\boldsymbol{u}\cdot\boldsymbol{c}_{t}=1$ for all $t\in[T]$, and finding such a $\boldsymbol{u}$ is typically impossible. When there does exist a $\boldsymbol{u}$ that satisfies this constraint, the reward relative to $\boldsymbol{u}$ is essentially the same as for us (Blum & Mansour, 2007). However, there will often be another value of $\boldsymbol{u}$ that will give us a much better bound, as we shall show in Section 6.
@@ 4 The CBA Algorithm
$\mathcal{V}_{t}:=\{\boldsymbol{v}\in\mathbb{R}_{+}^{E}\,|\,\|\boldsymbol{\pi}_{t}(\boldsymbol{v})\|_{1}\leq 1\}$
@@
Like mirror descent, CBA maintains, on each trial $t\in[T]$, a weight vector $\boldsymbol{w}_{t}\in\mathbb{R}_{+}^{E}$. However, unlike mirror descent on the simplex, we do not keep $\boldsymbol{w}_{t}$ normalized, but we will instead project it into $\mathcal{V}_{t}$ at the start of trial $t$, producing a vector $\tilde{\boldsymbol{w}}_{t}$. Also, unlike mirror descent, CBA does not use the actual gradient (which it does not know) of $\rho_{t}$ at $\tilde{\boldsymbol{w}}_{t}$, but uses an unbiased estimator. Specifically, on each trial $t\in[T]$ , CBA does the following:
@@
    Set $\tilde{\boldsymbol{w}}_{t}\leftarrow\operatorname{argmin}_{\boldsymbol{v}\in\mathcal{V}_{t}}\Delta(\boldsymbol{v},\boldsymbol{w}_{t})$
@@
    Randomly construct a vector $\boldsymbol{g}_{t}\in\mathbb{R}^{E}$ such that $\mathbb{E}[\boldsymbol{g}_{t}]=\nabla\rho_{t}(\tilde{\boldsymbol{w}}_{t})$
@@
    Set $\boldsymbol{w}_{t+1}\leftarrow\operatorname{argmin}_{\boldsymbol{v}\in\mathbb{R}_{+}^{E}}(\eta\boldsymbol{g}_{t}\cdot(\tilde{\boldsymbol{w}}_{t}-\boldsymbol{v})+\Delta(\boldsymbol{v},\tilde{\boldsymbol{w}}_{t}))$
@@
$\mathbb{E}[r_{t,a_{t}}]=\rho_{t}(\tilde{\boldsymbol{w}}_{t})$ (1)
which confirms that $\rho_{t}$ is indeed our objective function at trial $t$. Once $r_{t,a_{t}}$ is revealed to us we can go ahead and construct the gradient estimator $\boldsymbol{g}_{t}$. It is important for our analysis that we construct this estimator in a specific way. Inspired by Exp4 we first define a reward estimator $\hat{\boldsymbol{r}}_{t}$ such that for all $a\in[K]$ we have:
@@
Since this function is monotonic decreasing, $\lambda$ can be found by interval bisection.

### v2

@@ Abstract
We study the classic problem of prediction with expert advice under bandit feedback. Our model assumes that one action, corresponding to the learner’s abstention from play, has no reward or loss on every trial. We propose the confidence-rated bandits with abstentions (CBA) algorithm, which exploits this assumption to obtain reward bounds that can significantly improve those of the classical Exp4 algorithm. Our problem can be construed as the aggregation of confidence-rated predictors, with the learner having the option to abstain from play. We are the first to achieve bounds on the expected cumulative reward for general confidence-rated predictors. In the special case of specialists we achieve a novel reward bound, significantly improving the previous bounds of SpecialistExp (treating abstention as another action). We discuss how CBA can be applied to the problem of adversarial contextual bandits with the option of abstaining from selecting any action. We are able to leverage a wide range of inductive biases, outperforming previous approaches both theoretically and in preliminary experimental analysis. Additionally, we achieve a reduction in runtime from quadratic to almost linear in the number of contexts for the specific case of metric space contexts.
@@ 1 Introduction
We study the classic problem of prediction with expert advice under bandit feedback. The problem is structured as a sequence of trials. During each trial, each expert recommends a probability distribution over the set of possible actions. The learner then selects an action and observes and incurs the (potentially negative) reward associated with that action on that particular trial. In practical applications, errors often lead to severe consequences, and consistently making predictions is neither safe nor economically practical. For this reason, the abstention option has gained a lot of interest in the literature, both in the batch and online setting (Chow, 1957, 1970; Hendrickx et al., 2021; Cortes et al., 2018). Similarly to previous works, this paper is based on the assumption that one of the actions always has zero reward: such an action is equivalent to an abstention of the learner from play. Besides the rewards being bounded, we make no additional assumptions regarding how the rewards or expert predictions are generated. In this paper, we present an efficient algorithm CBA (Confidence-rated Bandits with Abstentions) which exploits the abstention action to get reward bounds that can be dramatically higher than those of Exp4 (Auer et al., 2002). In the worst case, our reward bound essentially matches that of Exp4 so that CBA can be seen as a strict improvement, since the time-complexities of the two algorithms are, up to a factor logarithmic in the time horizon, identical in the general case.
@@
Our problem can also be seen as that of aggregating confidence-rated predictors (Blum and Mansour, 2007; Gaillard et al., 2014; Luo and Schapire, 2015) when the learner has the option of abstaining from taking actions. When the problem is phrased in this way, at the start of each trial, each predictor recommends a probability distribution over the actions (which now may not include an action with zero reward) but with a confidence rating. A low confidence rating can mean that either the predictor thinks that all actions are bad (so that the learner should abstain) or simply does not know which action is the best. Previous works on confidence-rated experts measure the performance of their algorithm in terms of the sum of scaled per-trial rewards. In contrast to previous algorithms, our approach allows for the derivation of bounds on the expected cumulative reward of CBA.
This formulation enables us to extend our work to the problem of adversarial contextual bandits with the abstention option, which has not been studied before. Previous work has considered the abstention option in the standard (context-free) adversarial bandit setting or in stochastic settings (Cortes et al., 2018, 2020; Neu and Zhivotovskiy, 2020), but not in the contextual and adversarial case. Moreover, their results and methods cannot be applied to confidence-rated predictors. To get more intuition on this setup, we can think of any deterministic policy that maps contexts into actions. Any such policy can be viewed as a classifier, with foreground classes associated with each action and a background class associated with abstaining. Our learning bias is represented by a set of information we refer to as the basis, which we formally define later. It encodes contextual structural assumptions that hold exclusively for the foreground classes and are provided to the algorithm a priori. A particular type of basis is generated by a set of potential clusters that can overlap. Alternatively, a basis can also be created using balls generated by any kind of distance function, which groups contexts believed to be close together. For this latter family of basis, we can also achieve a significant speedup in the per-trial time complexity of CBA. This result is very different (and incomparable) to other results about adversarial bandits in metric spaces (Pasteris et al., 2023b, a).
@@ 1.1 Additional related work
The non-stochastic multi-armed bandit problem, initially introduced by Auer et al. (2002), has been a subject of significant research interest. Auer et al. (2002) also considered the multi-armed bandit problem with expert advice, introducing the Exp4 algorithm. Exp4 evolved the field of multi-armed bandits to encompass more complex scenarios, particularly the contextual bandit (Lattimore and Szepesvári, 2020). Contextual bandits are an extension of the classical multi-armed bandit framework, where an agent makes a sequence of decisions while taking into account contextual information. Our work is also related to the multi-class classification with bandit feedback, called weak reinforcement (Auer and Long, 1999). An action in our bandit setting corresponds to a class in the multi-class classification framework.
As discussed in the introduction, a key aspect of this work is the option to abstain from making any decision. In the batch setting (Chow, 1957, 1970), this option is usually referred to as “rejection”. These works study whether to use or reject a specific model prediction based on specific requests (see Hendrickx et al. (2021) for a survey). In online learning, “rejection” can be the possibility of abstention by the learner. These works usually rely on a cost associated with the abstention action.  Neu and Zhivotovskiy (2020) studied the magnitude of the cost associated with abstention in an expert setting with bounded losses. They state that if the cost is lower than half of the amplitude of the interval of the loss, it is possible to derive bounds that are independent of the time. In Cortes et al. (2018), a non-contextual and partial information setting with the option of abstention is studied. The sequel model (Cortes et al., 2020) regards this model as a special case of their stochastic feedback graph model. Schreuder and Chzhen (2021) studied the fairness setting when using the option of abstaining as it may lead to discriminatory predictions.
One specific scenario where prior algorithms can establish cumulative reward bounds is as follows: on any given trial, the predictors are specialists (Freund et al., 1997), having either full confidence (a.k.a. awake) or no confidence (a.k.a. asleep). The SpecialistExp algorithm by Herbster et al. (2021), a bandit version of the standard specialist algorithm, achieves regret bounds with respect to any subset of specialists where exactly one specialist is awake on each trial. We differ from this work as abstention is an algorithmic choice. Instead of sleeping in the rounds where the specialist is not active, the specialist will vote for abstention, which is a proper action of our algorithm. In Section 5.2, we present an illustrative problem involving learning balls in a space equipped with a metric. This example demonstrates our capability to significantly improve on SpecialistExp. For this problem, we also present subroutines that significantly speed up CBA.
@@ 2 Problem formulation and notation
    The reward incurred by action $a_{t}$ on trial $t$ (which is in $[-1,1]$) is revealed to the learner.
We note that the experts’ suggestions and the rewards (associated with each action) are chosen a-priori and hence do not depend on the learner’s actions. The aim of the learner is to maximize the cumulative reward obtained by its selected actions. As discussed in Section 1, we consider the case in which there is an action (the abstention action) that incurs zero reward on every trial.
@@
Any vector $\boldsymbol{s}\in\mathcal{A}$ represents the probability distribution over actions which assigns, for all $a\in[K]$, a probability of $s_{a}$ to action $a$, and assigns a probability of $1-\|\boldsymbol{s}\|_{1}$ to the abstention action $\square$, where $\|\boldsymbol{s}\|_{1}$ denotes 1-norm of $\boldsymbol{s}$. We write $a\sim\boldsymbol{s}$ to represent that action $a$ is drawn from the probability distribution $\boldsymbol{s}$. We will refer to the elements of the set $\mathcal{A}$ as stochastic actions.
A policy is any element of $\mathcal{A}^{T}$ (noting that any such policy is a matrix in $[0,1]^{T\times K}$). Any policy $\boldsymbol{e}\in\mathcal{A}^{T}$ defines a stochastic sequence of actions: on every trial $t\in[T]$ an action $a\in[K]\cup\{\square\}$ being drawn as $a\sim\boldsymbol{e}_{t}$. Note that if the learner plays according to a policy $\boldsymbol{e}\in\mathcal{A}^{T}$, then on each trial $t$ it obtains an expected reward of $\boldsymbol{r}_{t}\cdot\boldsymbol{e}_{t}$, where the operator $\cdot$ denotes the dot product. Note that each expert is equivalent to a policy. Thus, for all $i\in[E]$ we denote the $i$-th expert by $\boldsymbol{e}^{i}\in\mathcal{A}^{T}$. Hence, at the start of each trial $t\in[T]$, the learner views the sequence $\langle\boldsymbol{e}^{i}_{t}\,|\,i\in[E]\rangle$.
We can also view the experts as confidence-rated predictors over the set $[K]$: for each $i\in[E]$ and $t\in[T]$, the vector $\boldsymbol{e}^{i}_{t}$ can be viewed as suggesting the probability distribution $\boldsymbol{e}^{i}_{t}/\|\boldsymbol{e}^{i}_{t}\|_{1}$ over $[K]$, but with confidence $\|\boldsymbol{e}^{i}_{t}\|_{1}$. We denote this confidence by $c_{t,i}:=\|\boldsymbol{e}^{i}_{t}\|_{1}$ and write $\boldsymbol{c}_{t}:=(c_{t,1},\dots,c_{t,E})$.
In this work, we will refer to the unnormalized relative entropy defined by:
$\Delta(\boldsymbol{u},\boldsymbol{v}):=\sum_{i\in[E]}u_{i}\ln\left(\frac{u_{i}}{v_{i}}\right)-\|\boldsymbol{u}\|_{1}+\|\boldsymbol{v}\|_{1}$
@@ 3 Main result
Our main result is represented by a bound on the cumulative reward of our algorithm CBA. We note that any weight vector $\boldsymbol{u}\in\mathbb{R}_{+}^{E}$ induces a matrix $\boldsymbol{\pi}(\boldsymbol{u})\in\mathbb{R}_{+}^{T\times K}$ defined by
@@
which is the linear combination of the experts with coefficients given by $\boldsymbol{u}$. However, only some of such linear combinations generate valid policies. Thus, we define
$\mathcal{V}:=\{\boldsymbol{u}\in\mathbb{R}_{+}^{E}\,|\,\boldsymbol{\pi}(\boldsymbol{u})\in\mathcal{A}^{T}\}$
as the set of all weight vectors that generate valid policies. Particularly, note that $\boldsymbol{u}\in\mathcal{V}$ if and only if, on every trial $t$, the weighted sum of the confidences $\boldsymbol{u}\cdot\boldsymbol{c}_{t}$ is no greater than one. Given some $\boldsymbol{u}\in\mathcal{V}$, we define
@@
The following theorem (proved in Appendix A) allows us to bound the regret of CBA with respect to any valid linear combination $\boldsymbol{u}$ of experts.
@@
$\sum_{t\in[T]}\mathbb{E}[r_{t,a_{t}}]\geq\mathbb{E}[\rho(\boldsymbol{u})]-\frac{\Delta(\boldsymbol{u},\boldsymbol{w}_{1})}{\eta}-\eta(12K+2)T\,,$
where the expectations are with respect to the randomization of CBA’s strategy. The per-trial time complexity of CBA is in $\mathcal{O}(KE)$.
We now compare our bound to those of previous algorithms. Firstly, Exp4 can only achieve bounds relative to a $\boldsymbol{u}\in\mathcal{V}$ with $\|\boldsymbol{u}\|_{1}=1$ , in which case it essentially matches our bound but with $212K+$ 2 replaced by $88K+$ 8. Hence, for any $\boldsymbol{u}\in\mathcal{V}$ the Exp4 bound essentially replaces the term $\rho(\boldsymbol{u})$ in our bound by $\rho(\boldsymbol{u})/\|\boldsymbol{u}\|_{1}$. Note that $\|\boldsymbol{u}\|_{1}$ could be as high as the number of experts which implies we can dramatically outperform Exp4111Precisely, if for each expert there exists a trial in which the confidence is 1, then we have $0\leq\|u\|_{1}\leq E$. Otherwise can be high as $0\leq\|u\|_{1}\leq E/c^{*}$, where $c^{*}=\max_{t\in[T]}c_{t}^{i}$..
Secondly, when viewing our experts as confidence-rated predictors, we note that previous algorithms for this setting only give bounds on a weighted sum of the per-trial rewards where the weight on each trial is $\boldsymbol{u}\cdot\boldsymbol{c}_{t}$ for some $\boldsymbol{u}\in\mathcal{V}$. This is only a cumulative reward bound when $\boldsymbol{u}\cdot\boldsymbol{c}_{t}=1$ for all $t\in[T]$, and finding such a $\boldsymbol{u}$ is typically impossible. When there does exist $\boldsymbol{u}$ that satisfies this constraint, the reward relative to $\boldsymbol{u}$ is essentially the same as for us (Blum and Mansour, 2007). However, there will often be another value of $\boldsymbol{u}$ that will give us a much better bound, as we show in Section 5.2.
@@ 4 The CBA algorithm
$\mathcal{V}_{t}:=\{\boldsymbol{v}\in\mathbb{R}_{+}^{E}\,|\,\|\boldsymbol{\pi}_{t}(\boldsymbol{v})\|_{1}\leq 1\}\,,$
@@
Like mirror descent, CBA maintains, on each trial $t\in[T]$, a weight vector $\boldsymbol{w}_{t}\in\mathbb{R}_{+}^{E}$. However, unlike mirror descent on the simplex, we do not keep $\boldsymbol{w}_{t}$ normalized, but we will instead project it into $\mathcal{V}_{t}$ at the start of trial $t$, producing a vector $\tilde{\boldsymbol{w}}_{t}$. Also, unlike mirror descent, CBA does not use the actual gradient (which it does not know) of $\rho_{t}$ at $\tilde{\boldsymbol{w}}_{t}$, but (inspired by the Exp3 algorithm) uses an unbiased estimator instead. Specifically, on each trial $t\in[T]$ , CBA does the following:
@@
    Set $\tilde{\boldsymbol{w}}_{t}\leftarrow\operatorname{argmin}_{\boldsymbol{v}\in\mathcal{V}_{t}}\Delta(\boldsymbol{v},\boldsymbol{w}_{t})$.
@@
    Randomly construct a vector $\boldsymbol{g}_{t}\in\mathbb{R}^{E}$ such that $\mathbb{E}[\boldsymbol{g}_{t}]=\nabla\rho_{t}(\tilde{\boldsymbol{w}}_{t})$.
@@
    Set $\boldsymbol{w}_{t+1}\leftarrow\operatorname{argmin}_{\boldsymbol{v}\in\mathbb{R}_{+}^{E}}(\eta\boldsymbol{g}_{t}\cdot(\tilde{\boldsymbol{w}}_{t}-\boldsymbol{v})+\Delta(\boldsymbol{v},\tilde{\boldsymbol{w}}_{t}))$.
@@
$\mathbb{E}[r_{t,a_{t}}]=\rho_{t}(\tilde{\boldsymbol{w}}_{t})\,,$ (1)
which confirms that $\rho_{t}$ is our objective function at trial $t$. Once $r_{t,a_{t}}$ is revealed to us we can proceed to construct the gradient estimator $\boldsymbol{g}_{t}$. It is important that we construct this estimator in a specific way. Inspired by Exp4 we first define a reward estimator $\hat{\boldsymbol{r}}_{t}$ such that for all $a\in[K]$ we have:
@@
Since this function is monotonic decreasing, $\lambda$ can be found by interval bisection. For this computation step, we treat our numerical precision as a constant in our time complexity. In Appendix A.1, we show that, even if the numerical precision is unbounded, we incur a time complexity equal to that of Exp4, up to a factor logarithmic in $T$, adding only 1 to the regret.